I am having problems solving this word problem.
The quotient of a number and 3 is the same as the difference
of the number and two. Find the number
I came up with
x+3=y+2
I am not sure if this is a correct equation for this problem. I also need help solving it.
Thanks
Megan

Answer by nerdybill(5412)

Hi Megan,
.
If we:
Let x = "a number"
.
Then from: "quotient of a number and 3" we get:
x/3
.
And, from: "the difference of the number and two" we get:
x-2
.
In the problem, it said that the two terms above are the same giving us:
x/3 = x-2
.
Now, we solve for x:
x/3 = x-2
multiplying both sides by 3:
x = 3(x-2)
x = 3x - 6
subtracting x from both sides:
0 = 2x - 6
adding 6 to both sides:
6 = 2x
finally, dividing both sides by 2 gives us:
3 = x
.
Write back if you have questions,
Bill